# Gauge coupling jump and small instantons from a large non-minimal coupling

### v1

@@ Abstract
If a scalar field couples to the Ricci scalar with a large non-minimal coupling, the Standard Model coupling parameters can differ above and below a intermediate field range of the scalar due to the non-renormalizablity. In this paper, we study, for the first time, the threshold effects on a gauge coupling in both Metric and Palatini formulations of the gravity. We find that the gauge coupling naturally jumps around this intermediate scale since counter terms for the renormalization behave so. If the gauge coupling becomes strong with a large scalar field value due to this effect, there can be an enhanced small instanton contribution, the dilute gas approximation of which is justified because the gauge sector decouples when the scalar wave mode is very short. Using these findings, we discuss the QCD axion quality problem, the heavy QCD axion, the QCD axion abundance, and the suppression of isocurvature perturbations. We show that axion physics may differ substantially from naïve expectations when we introduce a large non-minimal coupling for any scalar field.
@@ 1 Introduction
The Higgs inflation [16, 17] (see also [18]), in which $\phi$ is the Higgs boson, was considered as a minimal possibility for inflation. A non-minimal coupling between the Higgs and Ricci scalar, $\xi$, is much larger than unity in this scenario. In this framework, the Standard Model (SM) Yukawa and Higgs quartic couplings can differ between the current universe and the inflationary period due to so-called threshold effects [19, 20, 21, 22, 23]. This effect can be understood as arising from the UV-completion likely present in the intermediate region $M_{\rm pl}/\xi<\phi<M_{\rm pl}/\sqrt{\xi}$, where unitarity is violated [24, 25, 26, 27]. While this primarily applies to the metric formulation, the Palatini formulation [28, 29] has a weaker perturbativity bound [30]. Still threshold effects are expected according to renormalization of the model. Due to this coupling change, the electroweak (EW) vacuum has been discussed as being absolutely stabilized, making Higgs inflation possible [19]. Apart from this coupling shift, the Standard Model parameters had been thought to disfavor minimal Higgs inflation because the Higgs quartic coupling turns negative at energy scales above $10^{9-12}\,{\rm\,GeV}$. However, it was recently shown that the EW vacuum can be made absolutely stable, and the inflation can be successful by taking into account the would-be hilltop of the Higgs potential and a very large non-minimal coupling[31]. [^1] One can have smaller non-minimal coupling if the quartic coupling turns negative around $10^{16}(10^{14}){\rm\,GeV}$ with the metric (Palatini) formulation by using the same mechanism [31, 32].
@@ 2.2 Threshold effect for the QCD gauge coupling
With a large Higgs non-minimal coupling, shifts in the couplings across intermediate regime are predicted [19, 20, 21, 22, 23]. This phenomenon has been studied in the context of Yukawa and Higgs quartic couplings [19]. Here, we estimate the impact of quark Yukawa interactions, particularly those involving the top quark, on the QCD gauge coupling jump through loop corrections (see Refs. [73, 74, 75] for expressions of various two-loop corrections).
@@
We can also have other terms for the gauge coupling jump in the presence of the large non-minimal coupling. However, they should behave similarly to have the chiral Standard Model in the limit $\varphi\to\infty$ (see e.g. [76]), the tree-level gauge coupling should approach a constant so that the derivatives of the induced corrections (e.g. from the QCD potential) to the $\varphi$ potential is vanishing. In the low field range, those “higher dimensional” terms should behave as the usual higher dimensional terms, and we have the usual perturbative effective theory connecting the Standard Model. Similar behavior can be found in a different type of threshold effect Eq. (38) and from threshold effects to the other couplings [19, 20, 21, 22, 23].
@@ 3.2 Case of the Higgs inflation and a quality problem for QCD axion
One possible solution to the quality problem pointed out so far is to assume a minimal flavor and CP violation. In other words, we consider $Y_{u},Y_{d}$ to be the complex spurion field for the breaking of the chiral symmetry. This property is satisfied at any order of the perturbation either in a small $h$ regime or a large $h$ regime. This is because the “higher dimensional terms” may also satisfy the symmetry (which is only spontaneously broken by the spurion fields). This scenario requires no CP phases for other coupling parameters. For instance, we can have the threshold effect contribution in the form $c(|h|^{2})H\bar{Q}Y_{u}u$. Our assumption says that $c$ does not have CP phase, although $Y_{u}$ may have. This setup still has a strong CP problem because $\arg\det Y_{u}\det Y_{d}$ is non-vanishing in general. Thus we will solve it by the QCD axion.
@@ 4 Conclusions
More generically, we may consider a scalar field with the non-minimal coupling completely decoupled from any quark in the Jordan frame. Then we go to the Einstein frame to have
${\mathcal{L}}_{\rm quark}\supset-\frac{h}{\Omega(\varphi)}\left(\bar{\psi}_{u}\cdot Y_{u}\cdot\psi_{u}+\bar{\psi}_{d}\cdot Y_{d}\cdot\psi_{d}\right),$ (37)
where $\varphi$ and $h$ are irrelevant. Then the loop diagram 1 with $\varphi$ replaced by $h$ gives a divergence. We have to introduce a counter term of the form
$\propto\frac{1}{\Omega^{2}(\varphi)}(\mathop{\rm tr}[Y_{u}Y_{u}^{\dagger}]+\mathop{\rm tr}[Y_{d}Y_{d}^{\dagger}])\mathop{\rm tr}GG.$ (38)
As before, we need to write the term in the original Lagrangian. One can see that the behavior of $\Omega^{-2}$ with $\Omega^{2}\equiv 1+\xi^{n/2}\left(\frac{\phi}{M_{\rm pl}}\right)^{n}$ is similar to $(\partial_{\varphi}F)^{2}$ with $n>0$. Then $g^{-2}+\delta g^{-2}\Omega^{-2}$ has the same asymptotic behavior as Eq. (15).

### v2

@@ Abstract
If a scalar field couples to the Ricci scalar with a large non-minimal coupling, the Standard Model coupling parameters can differ above and below an intermediate field range of the scalar due to the non-renormalizability. In this paper, we study, for the first time, the threshold effects on a gauge coupling in both Metric and Palatini formulations of gravity. We find that the gauge coupling naturally jumps around this intermediate scale since counter terms for the renormalization behave so. If the gauge coupling becomes strong with a large scalar field value due to this effect, there can be an enhanced small instanton contribution, the dilute gas approximation of which is justified because the gauge sector decouples when the scalar wave mode is very short. Using these findings, we discuss the QCD axion quality problem, the heavy QCD axion, the axion abundance, and the suppression of isocurvature perturbations. We show that axion physics may differ substantially from naïve expectations when we introduce a large non-minimal coupling for any scalar field.
@@ 1 Introduction
The Higgs inflation [16, 17] (see also [18]), in which $\phi$ is the Higgs boson, was considered as a minimal possibility for inflation. A non-minimal coupling between the Higgs and Ricci scalar, $\xi$, is much larger than unity in this scenario. In this framework, the Standard Model (SM) Yukawa and Higgs quartic couplings can differ between the current universe and the inflationary period due to so-called threshold effects [19, 20, 21, 22, 23, 24, 25]. This effect can be understood as arising from the UV-completion likely present in the intermediate region $M_{\rm pl}/\xi<\phi<M_{\rm pl}/\sqrt{\xi}$, where unitarity is violated [26, 27, 28, 29]. While this primarily applies to the metric formulation, the Palatini formulation [30, 31] has a weaker perturbativity bound [32]. Still threshold effects are expected according to renormalization of the model (even if the non-minimal coupling is not large the threshold effect may exist [25]). Due to this coupling change, the electroweak (EW) vacuum has been discussed as being absolutely stabilized, making Higgs inflation possible [19]. Apart from this coupling shift, the Standard Model parameters had been thought to disfavor minimal Higgs inflation because the Higgs quartic coupling turns negative at energy scales above $10^{9-12}\,{\rm\,GeV}$. However, it was recently shown that the EW vacuum can be made absolutely stable, and the inflation can be successful by taking into account the would-be hilltop of the Higgs potential and a very large non-minimal coupling[33]. [^1] One can have smaller non-minimal coupling if the quartic coupling turns negative around $10^{16}(10^{14}){\rm\,GeV}$ with the metric (Palatini) formulation by using the same mechanism [33, 34].
@@ 2.2 Threshold effect for the QCD gauge coupling
With a large Higgs non-minimal coupling, shifts in the couplings across intermediate regime are predicted [19, 20, 21, 22, 23, 24, 25]. This phenomenon has been studied in the context of Yukawa and Higgs quartic couplings [19]. Here, we estimate the impact of quark Yukawa interactions, particularly those involving the top quark, on the QCD gauge coupling jump through loop corrections (see Refs. [75, 76, 77] for expressions of various two-loop corrections).
@@
We can also have other terms for the gauge coupling jump in the presence of the large non-minimal coupling. However, they should behave similarly to have the chiral Standard Model in the limit $\varphi\to\infty$ (see e.g. [78]), the tree-level gauge coupling should approach a constant so that the derivatives of the induced corrections (e.g. from the QCD potential) to the $\varphi$ potential is vanishing. In the low field range, those “higher dimensional” terms should behave as the usual higher dimensional terms, and we have the usual perturbative effective theory connecting the Standard Model. Similar behavior can be found from threshold effects to the other couplings [19, 20, 21, 22, 23, 24, 25].
@@ 3.2 Case of the Higgs inflation and a quality problem for QCD axion
One possible solution to the quality problem pointed out so far is to assume a minimal flavor and CP violation. In other words, we consider $Y_{u},Y_{d}$ to be the complex spurion field for the breaking of the chiral symmetry. This property is satisfied at any order of the perturbation either in a small $h$ regime or a large $h$ regime. This is because the “higher dimensional terms” may also satisfy the symmetry (which is only spontaneously broken by the spurion fields). This scenario requires no CP phases for other coupling parameters. For instance, we can include the threshold effect contribution and write the Yukawa interaction in the form $c(|h|^{2})H\bar{Q}Y_{u}u$ [^8] We can also consider, e.g., ${\rm det}{(Y_{u})}\tilde{c}(|h|^{2})H\bar{Q}Y_{u}u$, which satisfies our assumption and is CP violating, in addition to the leading contribution $c(|h|^{2})H\bar{Q}Y_{u}u$. The shift to the strong CP phase is expect to be $\mathcal{O}(10^{-10})$ by assuming $\tilde{c}$ and $c$ is the same order at $h\to\infty$. Here assume that $c=1,\tilde{c}=0$ when $h\to 0$. This minimal flavor and CP violating scneairo may be a novel solution to the quality problem. A model building for it may be interesting. Our assumption says that $c$ does not have CP phase, although $Y_{u}$ may have. This setup still has a strong CP problem because $\arg\det Y_{u}\det Y_{d}$ is non-vanishing in general. Thus we will solve it by the QCD axion.
@@ 4 Conclusions
More generically, we may consider a scalar field with the non-minimal coupling completely decoupled from any quark in the Jordan frame. Even in this case, when we go to the Einstein frame, we have the coupling with the Standard Model sector. For instance, $\varphi$ couples to the (canonically normalized) Standard Model like Higgs boson since it has a mass term. Other than the derivative couplings to the Higgs boson, from the mass term of the Higgs boson there is a portal coupling $\partial_{\varphi\varphi}\Omega^{-2}m_{h}^{2}$. This contributes to the loop diagram for the gluon self-coupling and there is a divergence even in the dimensional regularization and we need a counter term. The thereshold effect naturally exists for the gauge coupling.